A uniform beam PQ of length 40m and weight 10N is supported at P and Q.it carried a load of 4N at a point 10m from p.what is the reaction at Q
1
Expert's answer
2021-07-15T10:51:31-0400

The equation of equilibrium of moments gives

"R_Q*(40\:\rm m)-Mg*(20\:\rm m)-4N*(10\:\rm m)=0"

Hence

"R_Q=\frac{\rm 10\:N*(20\:\rm m)-4\:N*(10\:\rm m)}{40\:\rm m}=4\: \rm N"